How many grams of Cu and what volume of O2 gas will be obtained if 18.4 ampere electric current is passed through the electrolytic cell of CuSO4for 1 hour and 42 minutes between graphite electrodes at 208 K temperature and 1 bar pressure?
(Cu=63.5u O=16u) How to calculate?
1
Expert's answer
2012-09-14T08:52:29-0400
m = k · Q,
Q = I · t.
k=Mw/F=Mw/96500

m=Mw/96500 * I *
t

m(Cu)=63.5/96500 *18,4*6120=74.099
g

m(O2)=32/96500*18.4*6120=37.34
g

PV=m/Mr*RT

V=(37.34/32)*8.31*208 /105=19.21 L